Determine the molecular formula
for Ibuprofen, a common
remedy. Analysis of ibuprofen
yields a molar mass of 206 g/mol
and a percent composition of
75.7% C, 8.80% H, and 15.5% O.
1
Expert's answer
2020-02-17T07:09:15-0500

N(C)=(206×0.757)/12=13

N(H)=(206×0.0880)=18

N(O)(206×0.1550)/16=2


Answer: C13H18O2